OpenStudy (anonymous):

The photography club has 28 members. There are 12 boys in the club. What is the ratio of boys to girls in simplest form? A. B. C. D.

OpenStudy (anonymous):

http://prntscr.com/9vaecj

OpenStudy (anonymous):

@jhonyy9

OpenStudy (calculusxy):

\(\huge\frac{\text{BOYS}}{\text{GIRLS}}\) we know the number of boys are 12 but we are not sure of the number of girls. however, one thing that we can use to find the # of boys is by using the info of the total of 28 members. \(\large 12 + g = 28\) where \(g\) = the # of girls solve for \(g\) please.

jhonyy9 (jhonyy9):

28-12=16 girls ratio boys girls 12/16 = 3/4

OpenStudy (anonymous):

is it D or A

OpenStudy (calculusxy):

none of those two options. @jhonyy9 basically gave you the answer.

OpenStudy (anonymous):

oh i did not see that